Calculus II, Fall 2014 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 14

1. Section 15.7: Problems 21, 31, 39, 43, 45.

2. Section 15.8: Problems 5, 9, 15, 23.

3. Section 16.1: Problems 15, 23, 25, 29.

4. Section 16.2: Problems 19, 23, 25, 27, 29, 35, 47.